# for the Baryon Scattering (BaSc) Collaboration

### v1

@@ Acknowledgements.
We thank Sinya Aoki for comments on the manuscript and results. The results presented here utilized: the Summit Supercomputer at Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725; the Lassen Supercompuer through the LLNL Multiprogrammatic and Institutional Computing program for Grand Challenge allocations; the Frontera Supercomputer Stanzione et al. (2020) at the Texas Advanced Computing Center (TACC); the Perlmutter Supercomputer at the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC awards NP-ERCAP0021598, NP-ERCAP0017492, and NP-ERCAP0015497. The chroma_laph and last_laph libraries were used to perform the sLapH calculations, making use of QDP++ USQCD (2020) and CHROMA Edwards and Joo (2005) software libraries. The quark propagator solves were performed with the QUDA library Clark et al. (2010); Babich et al. (2011); qud and the final contractions used contraction_optimizer Hörz (2019). The LALIBE library lal , branch feature/mp_nn was used to perform computations of the HAL QCD potential. Analysis code of the correlation functions is available in the git repository https://github.com/cosmon-collaboration/nn_c103_analysis which makes use of lsqfit Lepage (2020a), gvar Lepage (2020b) and TwoHadronsInBox Morningstar et al. (2017). This work was supported in part by the U.S. National Science Foundation (NSF) under awards PHY-1913158 and PHY-2209167 (C.M., S.S.), under award PHY-2209185 (A.S.), the NSF Faculty Early Career Development Program (CAREER) under award PHY-2047185 (A.N.), and by the Graduate Research Fellowship Program under Grant No. DGE-2040435 (J.M.). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF. This work was also supported in part by the U.S. Department of Energy (DOE), Office of Science, Office of Nuclear Physics, under grant contract numbers DE-AC02-05CH11231 (A.WL., J.M.), DE-SC0020250 (A.S.M.), DE-SC0004658 (K.M.), the DOE Topical Collaboration “Nuclear Theory for New Physics”, award No. DE-SC0023663 (A.N.,A.S.,A.WL,C.M.,J.M.), and U.S. DOE, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education (ORISE) for the DOE. ORISE is managed by ORAU under contract number DESC0014664 (J.M.). This work is supported by Lawrence Livermore National Security, LLC DE-AC52-07NA27344 (P.M.V., A.S.M.), Neutrino Theory Network Program Grant DE-AC02-07CHI11359 (A.S.M.). This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC05-00OR2272 (H.MC.). This work was supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through grant 513989149 (A.S.).
@@ Appendix A Correlator Analysis Systematics
Figure 27: Stability of the ground state interaction energy , ${\delta}_{00}$ for all 15 irreps/levels used in this work for the deuteron. Note that the two levels in $A_{2}(1)$ have overlapping values of ${\delta}_{00}$.
Figure 28: Stability of the ground state interaction energy , ${\delta}_{00}$ for all 8 irreps/levels used in this work in the di-neutron channel.
@@ B.1 Inelastic Excited State Contamination
$C_{\rm NN}^{el}(t)$ $=e^{-2M_{N}t}\sum_{n=0}\psi_{n}(\mathbf{r})Z^{E}_{n}e^{-{\Delta}^{\rm NN,E}_{n}t}\,,$
$C_{\rm NN}^{inel}(t)$ $=e^{-2M_{N}t}\sum_{n=1}\Psi_{n}(\mathbf{r})Z^{I}_{n}e^{-{\Delta}^{\rm NN,I}_{n}t}\,,$ (45)
where the energy gaps are ${\Delta}^{\rm NN,E}_{n}=2\sqrt{M_{N}^{2}+q_{n}^{2}}-2M_{N}$ and ${\Delta}^{\rm NN,I}_{n}=E^{\rm NN,I}_{n}-2M_{N}$. Inserting these expressions into the ratio correlation functions and retaining only the leading inelastic excited state contributions, one has
$R(t,\mathbf{r})$ $=\frac{C_{\rm NN}(t)}{C_{\rm N}(t)^{2}}=R^{E}(t,\mathbf{r})+R^{I}(t,\mathbf{r})$
@@
Under the approximating that the first inelastic state, $\Psi_{1}(\mathbf{r})$ is an asymptotic scattering state that does not couple to other states, it will satisfy similar relations with a different potential,
@@ B.3 Di-neutron potential analysis
Figure 30: We show the resulting potential and phase shift analysis as in Figs. 18, 19, 20, 22, 23, and 24, except for the di-neutron channel.

### v2

@@ Acknowledgements.
We thank Sinya Aoki and Jeremy Green for comments on the manuscript and results. The results presented here utilized: the Summit Supercomputer at Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725; the Lassen Supercompuer through the LLNL Multiprogrammatic and Institutional Computing program for Grand Challenge allocations; the Frontera Supercomputer Stanzione et al. (2020) at the Texas Advanced Computing Center (TACC); the Perlmutter Supercomputer at the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC awards NP-ERCAP0021598, NP-ERCAP0017492, and NP-ERCAP0015497. The chroma_laph and last_laph libraries were used to perform the sLapH calculations, making use of QDP++ USQCD (2020) and CHROMA Edwards and Joo (2005) software libraries. The quark propagator solves were performed with the QUDA library Clark et al. (2010); Babich et al. (2011); qud and the final contractions used contraction_optimizer Hörz (2019). The LALIBE library lal , branch feature/mp_nn was used to perform computations of the HAL QCD potential. The correlation function analysis code makes use of lsqfit Lepage (2020a), gvar Lepage (2020b) and TwoHadronsInBox Morningstar et al. (2017). This work was supported in part by the U.S. National Science Foundation (NSF) under awards PHY-1913158 and PHY-2209167 (C.M., S.S.), under award PHY-2209185 (A.S.), the NSF Faculty Early Career Development Program (CAREER) under award PHY-2047185 (A.N.), and by the Graduate Research Fellowship Program under Grant No. DGE-2040435 (J.M.). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF. This work was also supported in part by the U.S. Department of Energy (DOE), Office of Science, Office of Nuclear Physics, under grant contract numbers DE-AC02-05CH11231 (A.WL., J.M.), DE-SC0020250 (A.S.M.), DE-SC0004658 (K.M.), the DOE Topical Collaboration “Nuclear Theory for New Physics”, award No. DE-SC0023663 (A.N.,A.S.,A.WL,C.M.,J.M.), and U.S. DOE, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education (ORISE) for the DOE. ORISE is managed by ORAU under contract number DESC0014664 (J.M.). This work is supported by Lawrence Livermore National Security, LLC DE-AC52-07NA27344 (P.M.V., A.S.M.), Neutrino Theory Network Program Grant DE-AC02-07CHI11359 (A.S.M.). This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC05-00OR2272 (H.MC.). This work was supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through grant 513989149 (A.S.). This work was supported in part by ERC grant StrangeScatt-101088506 (J.B.).
@@ Appendix A Correlator Analysis Systematics
Refer to caption Refer to caption
(GEVP times, $t_{0}-t_{d}$) (blocking size, $N_{b}$)
Refer to caption Refer to caption
(the final NN time $t_{f}^{\rm NN}$) (the minimum N time $t_{\rm min}^{\rm N}$)
Figure 28: Stability of the ground state interaction energy , ${\delta}E_{00}$ for all 15 irreps/levels used in this work for the deuteron. Note that the two levels in $A_{2}(1)$ have overlapping values of ${\delta}E_{00}$.
Refer to caption Refer to caption
(GEVP times, $t_{0}-t_{d}$) (blocking size, $N_{b}$)
Refer to caption Refer to caption
(the final NN time $t_{f}^{\rm NN}$) (the minimum N time $t_{\rm min}^{\rm N}$)
Figure 29: Stability of the ground state interaction energy , ${\delta}E_{00}$ for all 8 irreps/levels used in this work in the di-neutron channel.
@@ B.1 Inelastic Excited State Contamination
$C_{\rm NN}^{el}(t,\textbf{r})$ $=e^{-2M_{N}t}\sum_{n=0}\psi_{n}(\mathbf{r})Z^{E}_{n}e^{-{\Delta}^{\rm NN,E}_{n}t}\,$,
$C_{\rm NN}^{inel}(t,\textbf{r})$ $=e^{-2M_{N}t}\sum_{n=1}\Psi_{n}(\mathbf{r})Z^{I}_{n}e^{-{\Delta}^{\rm NN,I}_{n}t}\,$, (50)
where the energy gaps are ${\Delta}^{\rm NN,E}_{n}=2\sqrt{M_{N}^{2}+\mathbf{q}_{n}^{2}}-2M_{N}$ and ${\Delta}^{\rm NN,I}_{n}=E^{\rm NN,I}_{n}-2M_{N}$. Inserting these expressions into the ratio correlation functions and retaining only the leading inelastic excited state contributions, one has
$R(t,\mathbf{r})$ $=\frac{C_{\rm NN}(t,\textbf{r})}{C_{\rm N}(t)^{2}}=R^{E}(t,\mathbf{r})+R^{I}(t,\mathbf{r})$
@@
Under the approximation that the first inelastic state, $\Psi_{1}(\mathbf{r})$ is an asymptotic scattering state that does not couple to other states, it will satisfy similar relations with a different potential,
@@ B.3 Di-neutron potential analysis
Figure 31: We show the resulting potential and phase shift analysis as in Figs. 19, 20, 21, 23, 24, and 25, except for the di-neutron channel.
B.4 Constructing the potential
The code we use to compute the HAL QCD potential, the LALIBE library lal , branch feature/mp_nn, produces NN correlation functions in relative momentum space with $\mathbf{P}_{\rm tot}=0$. As discussed in the main text, they were generated with zero-momentum wall sources for each nucleon.
B.4.1 $A_{1g}$ projection
We first need to eliminate all partial waves other than $\ell=0$. Most of them can be quickly removed by projection to $A_{1g}$, leaving only $\ell=0,4,8,10,\ldots$. We perform the projection on the momentum space correlation function in a two step process by averaging over all the symmetry transformations on the data that leave an element of $A_{1g}$ invariant:
def A1g_projection(NN):
# pseudo code
# positive parity projection
NN_g = NN( p_x, p_y, p_z)
NN_g += NN(-p_x, p_y, p_z)
NN_g += NN( p_x,-p_y, p_z)
NN_g += NN( p_x, p_y,-p_z)
NN_g += NN(-p_x,-p_y, p_z)
NN_g += NN(-p_x, p_y,-p_z)
NN_g += NN( p_x,-p_y,-p_z)
NN_g += NN(-p_x,-p_y,-p_z)
NN_g = NN_g / 8.0
# equivalent axes swap
# 0=p_x, 1=p_y, 2=p_z
NN_A1g = NN_g
NN_A1g += NN_g.swap(0,1)
NN_A1g += NN_g.swap(1,2)
NN_A1g += NN_g.swap(2,0)
NN_A1g += (NN_g.swap(0,1)).swap(1,2)
NN_A1g += (NN_g.swap(0,2)).swap(2,1)
NN_A1g = NN_A1g / 6.0
return NN_A1g
Later steps in the extraction of the HAL QCD potential require a position space representation Eq. (35) which is produced by a Fourier Transform back to relative coordinate space. This correlation function is then divided by the single nucleon one created with the same wall-source and projected to zero momentum at the sink, creating Eq. (37), which is then fed into Eq. (38) to solve for the central potential, Eq. (43).
B.4.2 Projection to $\ell=0$
We initially expected that $\ell=4,6,8,\ldots$ components of the $A_{1g}$ projected $R(t,r)$ would be small, but decided to perform a projection to $\ell=0$ to be sure.
[…]
A 1D interpolator is then constructed from the radial function samples and used to update the $R(t,r)$ values on the lattice sites. The update happens in a ball that does not reach the faces of the volume. Later steps will not use data in the corners of the volume.
This projection method is equivalent to the Misner method Misner (2004) used by the HAL QCD Collaboration Miyamoto et al. (2020).
The observed variation between the $\ell=0$ projected $R(t,r)$ and the $A_{1g}$ projected version it was derived from was less than $1\%$ and we observed no significant changes in the resulting phase shifts derived from the potential. Using the $\ell=0$ projection slightly reduces uncertainties from small changes in r that require large changes in direction to remain the lattice. Essentially the small $\ell=4$ contribution acts as noise. Fig. 32 shows the potential at $t=12$ with and without the $\ell=0$ projection.
Figure 32: The potential at $t=12$ without (left) and with (right) the $\ell=0$ projection described in the text.
B.4.3 Improved Time Derivatives
Last, we provide detail on the extraction of the time derivatives needed for Eq. (38). At each time and space point we sampled $R(t+\Delta t,r)$ with $\Delta t$ from $\left\{-2,-1,0,1,2\right\}$. This 5-element vector is called $\mathbf{v}$. We then fit a quartic polynomial with 5 coefficients $\mathbf{c}$ to reproduce $\mathbf{v}$. Let
$\mathbf{M}=\left(\begin{array}[]{rrrrr}1&{-2}&4&-8&16\\ 1&{-1}&1&-1&1\\ 1&0&0&0&0\\ 1&1&1&1&1\\ 1&2&4&8&16\end{array}\right).$ (64)
Each row of $\mathbf{M}$ contains powers of $\Delta t$ begining with 2-2 in the top row, eg. $M_{0i}=(-2)^{i}$, $M_{1i}=(-1)^{i}$, etc. Then
$\mathbf{M}\mathbf{c}=\mathbf{v}.$ (65)
Each row of $\Delta t$ powers is dotted with the as yet unknown coefficients $\mathbf{c}$ to produce the samples $R(t+\Delta t,r)$ in $\mathbf{v}$. Multiplication on both sides by $\mathbf{M}^{-1}$ produces the coefficients
$\mathbf{c}=\mathbf{M}^{-1}\mathbf{v}.$ (66)
The improved derivatives implemented in the code can be determined from derivatives of the polynomial evaluated at $\Delta t=0$.
$-\partial_{t}R(t)$ $=-c_{1}+\mathcal{O}\left(a^{4}\right)$ (67)
$=\frac{1}{12}\left[R(t+2)-R(t-2)\right]$
$\quad-\frac{2}{3}\left[R(t+1)-R(t-1)\right]+\mathcal{O}(a^{4})$
$\partial_{t}^{2}R(t)$ $=2c_{2}+\mathcal{O}\left(a^{4}\right)$
$=-\frac{1}{12}\left[R(t+2)+R(t-2)\right]-\frac{5R(t)}{2}$
$\quad+\frac{4}{3}\left[R(t+1)+R(t-1)\right]+\mathcal{O}\left(a^{4}\right)$
The spacing dependence of the error for the second derivative benefits from a symmetry induced cancellation, yielding the same order error as the first derivative.